OpenStudy (anonymous):

What is the equation of the line described below written in slope-intercept form? the line passing through point (0, 4), perpendicular to the line whose equation is y = -1/4x y = -4x y = -4x + 4 y = 1/4 x + 4 y = 4x + 4

OpenStudy (anonymous):

@satellite73

Mehek (mehek14):

perpendicular line's slope is the negative reciprocal

Mehek (mehek14):

so negative of -1/4 = 1/4

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

third?

Mehek (mehek14):

reciprocal of 1/4 = ?

OpenStudy (anonymous):

and reciprical is 4

OpenStudy (anonymous):

oh sorry ill leave this to you mehek

OpenStudy (anonymous):

4

Mehek (mehek14):

yes so D

OpenStudy (anonymous):

thank you

Mehek (mehek14):

yw :]